Falconmaster:

If D varies directly with t and D is 105 when t is 3, find D when t is 8. 35 280 840

Falconmaster:

b? #random guess

umm:

If it varies directly, then you can make a proportion. 105/3 = d/8 cross multiply (3)(d) = (105)(8) 3d = 840 D = 840/3 D = 280 <=== So you're correct.

Falconmaster:

YEAH BOI

umm:

LOL